OpenStudy (anonymous):

The gravitational force exerted on a baseball is 2.24 N down. A pitcher throws the ball horizontally with velocity 17.0 m/s by uniformly accelerating it along a straight horizontal line for a time interval of 183 ms. The ball starts from rest. (a) Through what distance does it move before its release? __________m (b) What are the magnitude and direction of the force the pitcher exerts on the ball? magnitude______N direction_______° above the horizontal

OpenStudy (anonymous):

d=vt since the acceleration is uniform just convert the 183 ms to s t=.183s. multiply the two to get 3.111meters

OpenStudy (anonymous):

v=vo + at^2 to get a in the x direction 17 - 0/(.183^2) = a = 507.629m/s^2 which seems really fast but also its over not even a fifth of a second that it is increasing. you can find the mass of the ball by dividing 2.24N down by 9.8 = .2285kg. F=ma=(.2285*507.629) = 116.03N in x direction. inverse sin of (2.24/116.03) = 1.106 degrees that the pitcher needs to throw up to compensate for the gravity down on the ball

OpenStudy (anonymous):

its all wrong =/

OpenStudy (anonymous):

it says its within 10% of the right answer though